What volume of CO2 would be formed if 5.00 L of propane burns, assuming that all of the gases are under the same conditions.


1
Expert's answer
2021-12-15T10:17:41-0500

C3H8 (g) + 5 O2 (g) -->3 CO2 (g) + 4 H2O

Molar mass of C3H8=44.1

5000/44.1=113.38

Molar mass of CO2= 44.01

= 44.01×113.38×3=14969.56ml